OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)]

OpenStudy (bahrom7893):

g(-2) = 6(-2)-1=-12-1=-13

OpenStudy (bahrom7893):

f[g(-2)]=f[-13] = -2*(-13)-1

OpenStudy (anonymous):

thank you i had it right but i changed it thinking i subtracted wrong i have another can you help?

OpenStudy (bahrom7893):

Sure, go ahead

OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4).

OpenStudy (bahrom7893):

f(x)-g(x) = (2x-2)-(3x-1) After that plug in x=4

OpenStudy (anonymous):

7 puto http://t.qkme.me/3ru3jt.jpg

OpenStudy (anonymous):

i did it twice and got different answers i got 17 or 5

OpenStudy (bahrom7893):

2x-3x-2+1=-x-1

OpenStudy (bahrom7893):

(f-g) = -x-1 (f-g)(4) = -4-1 = -5

OpenStudy (anonymous):

i was going with 5 just forgot the negative sign thanks so much!